# From Similarity to Superiority: Channel Clustering for Time Series Forecasting

### v1

@@ 4.2 CCM: Channel Clustering Module
Channel Clustering. Motivated by the above observations, we first initialize a set of $K$ cluster embedding $\{c_{1},\cdots,c_{K}\}$, where $c_{k}\in\mathbb{R}^{d}$, $d$ is the hidden dimension and $K$ is a hyperparameter. Given a multivariate time series $X\in\mathbb{R}^{T\times C}$, each channel in the input $X_{i}$ is transformed into a $d$-dimensional channel embedding $h_{i}$ through an MLP. The probability that a given channel $X_{i}$ is associated with the $k$-th cluster is the normalized inner-product of the cluster embedding $c_{k}$ and the channel embedding $h_{i}$, which is computed as
$p_{i,k}=\text{Normalize}(\frac{c_{k}^{\top}h_{i}}{\|c_{k}\|\|h_{i}\|})\in[0,1].\vspace{-0.1cm}$ (2)
The normalization operator ensures that $\sum_{k}p_{i,k}=1$ and validates the clustering probability distribution across $k$ clusters. We utilize reparameterization trick (Jang et al., 2016) to obtain the clustering membership matrix $\mathbf{M}\in\mathbb{R}^{C\times K}$ where $\mathbf{M}_{ik}\approx\text{Bernoulli}(p_{i,k})$. Higher probability $p_{i,k}$ results in $\mathbf{M}_{ik}$ close to 1, leading to deterministic existence of certain channel in the corresponding cluster.
@@
Cluster Loss. We further introduce a specifically designed loss function for the clustering quality, termed ClusterLoss, which incorporates both the alignment of channels with their respective clusters and the distinctness between different clusters in a self-supervised learning context. Let $\mathbf{S}\in\mathbb{R}^{C\times C}$ denote the channel similarity matrix with $\mathbf{S}_{ij}=\textsc{Sim}(X_{i},X_{j})$ defined in Eq. 1. The ClusterLoss is formulated as:
@@
Cluster-aware Feed Forward. Instead of using individual Feed Forward per channel in a CI manner or sharing one Feed Forward across all channels in a CD manner, we assign a separate Feed Forward to each cluster to capture the underlying shared time series patterns within the clusters. In this way, we use cluster identity to replace channel identity. Each Feed Forward is parameterized with a single linear layer due to its efficacy in time series forecasting (Zeng et al., 2022; Li et al., 2023a; Chen et al., 2023). Let $h_{\theta_{k}}(\cdot)$ represent the linear layer for the $k$-th cluster with weights $\theta_{k}$. $Z_{i}$ represents the hidden embedding of the $i$-th channel before the last layer. The final forecast is thereby averaged across the outputs of all cluster-aware Feed Forward with $\{p_{i,k}\}$ as weights, e.g., $Y_{i}=\sum_{k}p_{i,k}h_{\theta_{k}}(Z_{i})$ for the $i$-th channel. For computational efficiency, it is equivalent to $Y_{i}=h_{\theta^{i}}(Z_{i})$ with averaged weights $\theta^{i}=\sum_{k}p_{i,k}\theta_{k}$.
@@
Zero-shot Forecasting. The prototype embeddings acquired during the training phase serve as a compact representation of the pre-trained knowledge and can be harnessed for seamless knowledge transfer to unseen samples or new channels. We apply this knowledge to unseen instances by computing the clustering probability distribution on the pre-trained clusters, following Eq. 2. The clustering probability is subsequently used for averaging cluster-aware Feed Forward. We disable the cross-attention operation to fix the prototype embeddings in zero-shot forecasting scenarios. It is worth noting that zero-shot forecasting is adaptable to both univariate and multivariate scenarios. We refer to Appendix B for detailed discussion.
@@ 4.3 Complexity Analysis
CCM effectively strikes a balance between the CI and CD strategies. On originally CI models, CCM introduces strategic clustering on channels, which not only reduces the model complexity but also enhances their generalizability. Simultaneously, CCM increases the model complexity on originally CD models with negligible overhead for higher capacities. We refer to Figure 5 in Appendix C.5 for empirical analysis. Theoretically, the computational complexity of clustering probability computation (Eq. 2) and the cross-attention (Eq. 3) are $\mathcal{O}(KCd)$, where $K,C$ are the number of clusters and channels, respectively, and $d$ is the hidden dimension. One may also use more attention mechanisms (Kitaev et al., 2020; Wang et al., 2020; Shen et al., 2021) for efficiency. The complexity of cluster-aware Feed Forward scales linearly in $C,K$, and the forecasting horizon $H$.
@@ 5.1 Experimental Setup
Datasets. For long-term forecasting, we experiment on 9 popular benchmarking datasets across diverse domains (Zhou et al., 2021; Wu et al., 2021; Lai et al., 2018), including weather, traffic and electricity. M4 dataset (Makridakis et al., 2018) is used in short-term forecasting, which is a univariate dataset that covers time series across diverse domains and various sampling frequencies from hourly to yearly. We further provide a new stock time series dataset with 1390 univariate time series. Each time series records the price history of an individual stock spanning 10 years. Due to the potential significant fluctuations in stock performance across different companies, this dataset poses challenges for models capturing diverse and evolving stock patterns in financial markets. The statistics of long-term and short-term datasets are shown in Table 3 and Table 3.
Table 2: The statistics of datasets
in long-term forecasting. Horizon
is $\{96,192,336,720\}$.
@@
Stock (Ours) 10000 7/24
We follow standard protocols (Zhou et al., 2021; Wu et al., 2021, 2022) for data splitting on public benchmarking datasets. As for the stock dataset, we divide the set of stocks into train/validation/test sets with a ratio of 7:2:1. Therefore, validation/test sets present unseen samples (i.e., stocks) for model evaluation. This evaluation setting emphasizes the data efficiency aspect of time series models for scenarios where historical data is limited or insufficient for retraining from scratch given unseen instances. More details on datasets are provided in Appendix C.1.
Base Models and Experimental Details. CCM is a model-agnostic channel strategy that can be applied to arbitrary time series forecasting models for improved performance. We meticulously select four recent state-of-the-art time series models as base models: TSMixer (Chen et al., 2023), DLinear (Zeng et al., 2022), PatchTST (Nie et al., 2022) and TimesNet (Wu et al., 2022), which mainly cover three mainstream paradigms, including linear models, transformer-based and convolutional models. For fair evaluation, we use the optimal experiment configuration as provided in the official code to implement both base models and the enhanced version with CCM. All the experiments are implemented by PyTorch on a single NVIDIA RTX A6000 48GB GPU. Experiment configurations and implementations are detailed in Appendix C.3. Experimental results in the following sections are averaged on five runs with different random seeds. Refer to Appendix C.6 for standard deviation results.
@@ 5.2 Long-term Forecasting Results
We report the mean squared error (MSE) and mean absolute error (MAE) on nine real-world datasets for long-term forecasting evaluation in Table 3. The forecasting horizon is $\{96,192,336,720\}$. From the table, we observe that the model enhanced with CCM outperforms the base model in general. Specifically, CCM improves the long-term forecasting performance in $90.27\%$ cases in MSE and $84.03\%$ cases in MAE across 144 different experiment settings. Remarkably, CCM achieves a substantial boost on DLinear, with a significant reduction on MSE by $5.12\%$ and MAE by $3.04\%$. The last column of the table quantifies the average percentage improvement in terms of MSE/MAE, which underscores the consistent enhancement brought by CCM across all forecasting horizons and datasets.
Comparison with Regularization Solution. PRReg (Han et al., 2023) recently displays to improve the performance of CD models, by predicting residuals with a regularization term in the training objective. We evaluate the effectiveness of PRReg and our proposed CCM on long-term forecasting performance enhancement based on CI and CD models. Following the previous training setting (Han et al., 2023), we develop CI and CD versions for Linear (Zeng et al., 2022) and Transformer (Vaswani et al., 2017) and report MSE loss. Table 5 shows that CCM surpasses PRReg in most cases, highlighting its efficacy compared with regularization solutions. See full results in Appendix C.4.
@@ 5.3 Short-term Forecasting Results
Table 6: Short-term forecasting results on M4 dataset in terms of SMAPE, MASE, and OWA, the lower the better. The better performance in each setting is shown in bold.
@@
Table 7: Short-term forecasting results on stock dataset in terms of MAE and MSE, the lower the better. The forecasting horizon is $\{7,24\}$. The better performance in each setting is shown in bold.
Horizon Metric TSMixer + CCM DLinear + CCM PatchTST + CCM TimesNet + CCM IMP(%)
7 MSE 0.939 0.938 0.992 0.883 0.896 0.892 0.930 0.915 3.288
@@
In both M4 and stock datasets, we follow the univariate forecasting setting. For M4 benchmarks, we adopt the evaluation setting in prior works (Oreshkin et al., 2019) and report the symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA) in Table 6. As for the stock dataset, we implement MAE and MSE as metrics in Table 7. See more details on metrics in Appendix C.2.
Remarkably, the efficacy of CCM is consistent across all M4 sub-datasets with different sampling frequencies. Specifically, CCM outperforms the state-of-the-art linear model (DLinear) by a significant margin of $11.62\%$, and outperforms the best convolutional method TimesNet by $8.88\%$. In table 7, we also observe a significant performance improvement on the stock dataset, achieved by applying CCM. In the stock dataset, we test on new samples (i.e., univariate stock time series) that the model has not seen in the training phase to evaluate the model’s generalization and robustness. By memorizing cluster-specific knowledge from analogous samples, the model potentially captures various market trends and behaviors and thereby makes more accurate and informed forecasting.
5.4 Zero-shot Forecasting Results
@@
Existing time series models tend to be rigidly tailored to specific domains, leading to poor generalization on data from unseen domains. In contrast, CCM leverages learned prototypes to capture cluster-specific knowledge. This enables meaningful comparisons between unseen time series and pre-trained knowledge, facilitating accurate zero-shot forecasting. Following prior work (Jin et al., 2023), we adopt ETT collections (Zhou et al., 2021), where ETTh1 and ETTh2 are hourly recorded, while ETTm1 and ETTm2 are minutely recorded. “1” and “2” indicate two different regions where the datasets originated.
Table 8 shows MSE and MAE results on test datasets. CCM consistently improves the zero-shot forecasting capacity of base time series models in 48 scenarios, including generalization to different domains and different granularities. Specifically, based on the results, we make the following observations. (1) CCM exhibits more significant performance improvement with longer forecasting horizons, highlighting the efficacy of memorizing and leveraging pre-trained knowledge in zero-shot forecasting scenarios. (2) Achieving cross-granularity (i.e., different sampling frequencies) generalization presents greater challenges than cross-domain (i.e., different sampling regions) generalization. However, the result comparison between task 111 and 222 (or task 444 and 555 ) shows that CCM exhibits superiority in more challenging cross-granularity generalization scenarios. (3) CCM demonstrates a better effect on originally CI base models. For instance, the averaged improvement rates on two CI models, i.e., DLinear and PatchTST, are $10.48\%$ and $11.13\%$ respectively, while the improvement rates on TSMixer and TimesNet are $5.14\%$ and $9.63\%$. Overall, the experimental results verify the superiority and efficacy of CCM in enhancing zero-shot transfer forecasting.
@@ 5.5 Qualitative Visualization
Channel Clustering Visualization. Figure 2 presents the t-SNE visualization of channel and prototype embeddings within ETTh1 and ETTh2 datasets with DLinear as the base model. Each point represents a channel within a sample, with varying colors indicating different channels. In ETTh1 dataset, we discern a pronounced clustering of channels 0, 2, and 4, suggesting that they may be capturing related or redundant information within the dataset. Concurrently, channels 1, 3, 5, and 6 coalesce into another cluster. The similarity matrix in the lower left further corroborates these findings. Clustering is also observable in ETTh2 dataset, particularly among channels 0, 4, and 5, as well as channels 2, 3, and 6. Comparatively, channel 1 shows a dispersion among clusters, partly due to its capturing of unique or diverse aspects of the data that do not closely align with the features represented by any clusters. The clustering results demonstrate that CCM not only elucidates the intricate relationships and potential redundancies among the channels but also offers critical insights for feature analysis and enhancing the interpretability of time series models.
Figure 2: t-SNE visualization of channel and prototype embedding given by DLinear with CCM on (a) ETTh1 and (b) ETTh2 dataset. The lower left corner shows the similarity matrix between channels.
Weight Visualization of Cluster-aware Projection. Figure 3 depicts the weights visualization for the cluster-aware Feed Forward on ETTh1 and ETTm1 datasets, revealing distinct patterns that are indicative of the model’s learned features (Li et al., 2023a; Zeng et al., 2022; Yi et al., 2023). For instance, in the ETTm1 dataset, Cluster 0 shows bright diagonal striping patterns, which may suggest that it is primarily responsible for capturing the most dominant periodic signals in the corresponding cluster. In contrast, Cluster 1 exhibits denser stripes, indicating its role in refining the representation by capturing more subtle or complex periodicities that the first layer does not. The visualization implies the model’s ability to identify and represent periodicity in diverse patterns, which is crucial for time-series forecasting tasks that are characterized by intricate cyclic behaviors.
@@
5.6 Ablation Study on Cluster Number

### v2

@@ 4.2 CCM: Channel Clustering Module
Channel Clustering. Motivated by the above observations, we first initialize a set of $K$ cluster embeddings $\{c_{1},\cdots,c_{K}\}$, where $c_{k}\in\mathbb{R}^{d}$, $d$ is the hidden dimension and $K$ is a hyperparameter. Given a multivariate time series $X\in\mathbb{R}^{T\times C}$, each channel in the input $X_{i}$ is transformed into a $d$-dimensional channel embedding $h_{i}$ through an MLP. The probability that a given channel $X_{i}$ is associated with the $k$-th cluster is the normalized inner-product of the cluster embedding $c_{k}$ and the channel embedding $h_{i}$, which is computed as
$p_{i,k}=\text{Normalize}(\frac{c_{k}^{\top}h_{i}}{\|c_{k}\|\|h_{i}\|})\in[0,1].$ (2)
The normalization operator ensures that $\sum_{k}p_{i,k}=1$ and validates the clustering probability distribution across $k$ clusters. We utilize reparameterization trick [67] to obtain the clustering membership matrix $\mathbf{M}\in\mathbb{R}^{C\times K}$ where $\mathbf{M}_{ik}\approx\text{Bernoulli}(p_{i,k})$. Higher probability $p_{i,k}$ results in $\mathbf{M}_{ik}$ close to 1, leading to the deterministic existence of certain channels in the corresponding cluster.
@@
Cluster Loss. We further introduce a specifically designed loss function for the clustering quality, termed ClusterLoss, which incorporates both the alignment of channels with respective clusters and the distinctness between different clusters in a self-supervised context. Let $\mathbf{S}\in\mathbb{R}^{C\times C}$ denote the channel similarity matrix $\mathbf{S}_{ij}=\textsc{Sim}(X_{i},X_{j})$ defined in Eq. 1. The ClusterLoss is formulated as:
@@
Cluster-aware Feed Forward. Instead of using individual Feed Forward per channel in a CI manner or sharing one Feed Forward across all channels in a CD manner, we assign a separate Feed Forward to each cluster to capture the underlying shared time series patterns within the clusters. In this way, we use cluster identity to replace channel identity. Each Feed Forward is parameterized with a single linear layer due to its efficacy in time series forecasting [8, 15, 7]. Let $h_{\theta_{k}}(\cdot)$ represent the linear layer for the $k$-th cluster with weights $\theta_{k}$. $Z_{i}$ represents the hidden embedding of the $i$-th channel before the last layer. The final forecast is thereby averaged across the outputs of all cluster-aware Feed Forward with $\{p_{i,k}\}$ as weights, e.g., $Y_{i}=\sum_{k}p_{i,k}h_{\theta_{k}}(Z_{i})$ for the $i$-th channel. For computational efficiency, it is equivalent to $Y_{i}=h_{\theta^{i}}(Z_{i})$ with averaged weights $\theta^{i}=\sum_{k}p_{i,k}\theta_{k}$.
@@
Zero-shot Forecasting. Zero-shot forecasting is useful in time series applications where data privacy concerns restrict the feasibility of training models from scratch for unseen samples. The prototype embeddings acquired during the training phase serve as a compact representation of the pre-trained knowledge and can be harnessed for seamless knowledge transfer to unseen samples or new channels in a zero-shot setting. The pre-trained knowledge is applied to unseen instances by computing the clustering probability distribution on the pre-trained clusters, following Eq. 2, which is subsequently used for averaging cluster-aware Feed Forward. The cross-attention is disabled to fix the prototype embeddings in zero-shot forecasting. It is worth noting that zero-shot forecasting is applicable to both univariate and multivariate scenarios. We refer to Appendix B for detailed discussion.
@@ 4.3 Complexity Analysis
CCM effectively strikes a balance between the CI and CD strategies. On originally CI models, CCM introduces strategic clustering on channels, which not only reduces the model complexity but also enhances their generalizability. Simultaneously, CCM increases the model complexity on originally CD models with negligible overhead for higher capacities. We refer to Figure 5 for empirical analysis. Theoretically, the computational complexity of clustering probability computation (Eq. 2) and the cross-attention (Eq. 3) are $\mathcal{O}(KCd)$, where $K,C$ are the number of clusters and channels, respectively, and $d$ is the hidden dimension. One may also use other attention mechanisms [68, 69, 70] for efficiency. The complexity of cluster-aware Feed Forward scales linearly in $C,K$, and the forecasting horizon $H$.
@@ 5.1 Experimental Setup
Datasets. For long-term forecasting, we experiment on 9 popular benchmarking datasets across diverse domains [11, 12, 71], including weather, traffic and electricity. M4 dataset [72] is used in short-term forecasting, which is a univariate dataset that covers time series across diverse domains and various sampling frequencies from hourly to yearly. We further provide a new stock time series dataset with 1390 univariate time series. Each time series records the price history of an individual stock spanning 10 years. Due to the potential significant fluctuations in stock performance across different companies, this dataset poses challenges for capturing diverse and evolving stock patterns in financial markets. The statistics of long- and short-term datasets are shown in Table 3 and Table 3.
Table 2: The statistics of datasets in long-term forecasting. Horizon is $\{96,192,336,720\}$.
@@
Stock (New) 10000 7/24
We follow standard protocols [11, 12, 13] for data splitting on public benchmarking datasets. As for the stock dataset, we divide the set of stocks into train/validation/test sets with a ratio of 7:2:1. Therefore, validation/test sets present unseen samples (i.e., stocks) for model evaluation. This evaluation setting emphasizes the data efficiency aspect of time series models for scenarios where historical data is limited or insufficient for retraining from scratch given unseen instances. More details on datasets are provided in Appendix C.1.
Base Models and Experimental Details. CCM is a model-agnostic channel strategy that can be applied to arbitrary time series forecasting models for improved performance. We meticulously select four recent state-of-the-art time series models as base models: TSMixer [7], DLinear [8], PatchTST [21] and TimesNet [13], which mainly cover three mainstream paradigms, including linear models, transformer-based and convolutional models. For fair evaluation, we use the optimal experiment configuration as provided in the official code to implement both base models and the enhanced version with CCM. All the experiments are implemented with PyTorch on a single NVIDIA RTX A6000 48GB GPU. Experiment configurations and implementations are detailed in Appendix C.3. Experimental results in the following sections are averaged on five runs with different random seeds. Refer to Appendix C.6 for standard deviation results.
@@ 5.2 Long-term Forecasting Results
We report the mean squared error (MSE) and mean absolute error (MAE) on nine real-world datasets for long-term forecasting evaluation in Table 3. The forecasting horizon is $\{96,192,336,720\}$. From the table, we observe that the model enhanced with CCM outperforms the base model in general. Specifically, CCM improves long-term forecasting performance in $90.27\%$ cases in MSE and $84.03\%$ cases in MAE across 144 different experiment settings. Remarkably, CCM achieves a substantial boost on DLinear, with a significant reduction on MSE by $5.12\%$ and MAE by $3.04\%$. The last column of the table quantifies the average percentage improvement in terms of MSE/MAE, which underscores the consistent enhancement brought by CCM across all forecasting horizons and datasets. Intuitively, the CCM method is more useful in scenarios where channel interactions are complex and significant, which is usually the case in real-world data. See more analysis in Appendix C.5.
Comparison with Regularization Solution. PRReg [17] recently displays to improve the performance of CD models, by predicting residuals with a regularization term in the training objective. We evaluate the effectiveness of PRReg and our proposed CCM on long-term forecasting performance enhancement based on CI and CD models. Following the previous training setting [17], we develop CI and CD versions for Linear [8] and Transformer [41] and report MSE loss. Table 5 shows that CCM surpasses PRReg in most cases, highlighting its efficacy compared with regularization solutions. See full results in Appendix C.4.
@@ 5.3 Short-term Forecasting Results
In both M4 and stock datasets, we follow the univariate forecasting setting. For M4 benchmarks, we adopt the evaluation setting in prior works [42] and report the symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA). As for the stock dataset, we implement MAE and MSE as metrics in Table 6. See more details on metrics in Appendix C.2. Remarkably, the efficacy of CCM is consistent across all M4 sub-datasets with different sampling frequencies. Specifically, CCM outperforms the state-of-the-art linear model (DLinear) by a significant margin of $11.62\%$, and outperforms the best convolutional method TimesNet by $8.88\%$. In Table 6, we also observe a significant performance improvement on the stock dataset, achieved by applying CCM. In the stock dataset, we test on new samples (i.e., univariate stock time series) that the model has not seen during training to evaluate the model’s generalization and robustness. By memorizing cluster-specific knowledge from analogous samples, the model potentially captures various market trends and behaviors and thereby makes more accurate and informed forecasting.
Table 6: Short-term forecasting results on M4 dataset in terms of SMAPE, MASE, and OWA, and stock dataset in terms of MSE and MAE. The lower the better. The forecasting horizon is $\{7,24\}$ for the stock dataset. The better performance in each setting is shown in bold.
@@
Stock (Horizon 7) MSE 0.939 0.938 0.992 0.883 0.896 0.892 0.930 0.915 3.288
@@
5.4 Zero-shot Forecasting Results
Existing time series models tend to be rigidly tailored to a specific dataset, leading to poor generalization on unseen data. In contrast, CCM leverages learned prototypes to capture cluster-specific knowledge. This enables meaningful comparisons between unseen time series and pre-trained knowledge, facilitating accurate zero-shot forecasting. Following prior work [73], we adopt ETT collections [11], where ETTh1 and ETTh2 are hourly recorded, while ETTm1 and ETTm2 are minutely recorded. "1" and "2" indicate two different regions where the datasets originated. Table 7 shows MSE and MAE results on test datasets. CCM consistently improves the zero-shot forecasting capacity of base time series models in 48 scenarios, including generalization to different regions and different granularities. Specifically, based on the results, we make the following observations. (1) CCM exhibits more significant performance improvement with longer forecasting horizons, highlighting the efficacy of memorizing and leveraging pre-trained knowledge in zero-shot forecasting scenarios. (2) CCM demonstrates a better effect on originally CI base models. For instance, the averaged improvement rates on two CI models, i.e., DLinear and PatchTST, are $10.48\%$ and $11.13\%$ respectively, while the improvement rates on TSMixer and TimesNet are $5.14\%$ and $9.63\%$. Overall, the experimental results verify the superiority and efficacy of CCM in enhancing zero-shot forecasting and the practical value of generalization within closely related domains under varying conditions.
@@ 5.5 Qualitative Visualization
Channel Clustering Visualization. Figure 2 presents the t-SNE visualization of channel and prototype embeddings within ETTh1 and ETTh2 datasets with DLinear as the base model. Each point represents a channel within a sample, with varying colors indicating different channels. In ETTh1 dataset, we discern a pronounced clustering of channels 0, 2, and 4, suggesting that they may be capturing related or redundant information within the dataset. Concurrently, channels 1, 3, 5, and 6 coalesce into another cluster.
Figure 2: t-SNE visualization of channel and prototype embedding by DLinear with CCM on (a) ETTh1 and (b) ETTh2 dataset. The lower left corner shows the similarity matrix between channels.
The similarity matrix in the lower left further corroborates these findings. Clustering is also observable in ETTh2 dataset, particularly among channels 0, 4, and 5, as well as channels 2, 3, and 6. Comparatively, channel 1 shows a dispersion among clusters, partly due to its capturing of unique or diverse aspects of the data that do not closely align with the features represented by any clusters. The clustering results demonstrate that CCM not only elucidates the intricate relationships and potential redundancies among the channels but also offers critical insights for feature analysis and enhancing the interpretability of time series models.
Weight Visualization of Cluster-aware Projection. Figure 3 depicts the weights visualization for the cluster-aware Feed Forward on ETTh1 and ETTm1 datasets, revealing distinct patterns that are indicative of the model’s learned features [15, 8, 51].
@@
For instance, in the ETTm1 dataset, Cluster 0 shows bright diagonal striping patterns, which may suggest that it is primarily responsible for capturing the most dominant periodic signals in the corresponding cluster. In contrast, Cluster 1 exhibits denser stripes, indicating its role in refining the representation by capturing more subtle or complex periodicities that the first layer does not. The visualization implies the model’s ability to identify and represent periodicity in diverse patterns, which is crucial for time-series forecasting tasks that are characterized by intricate cyclic behaviors.
5.6 Ablation Studies